OpenStudy (anonymous):

A boat traveled 64 miles downstream and back. The trip downstream took 4 hours. The trip back (upstream) took 8 hours. Find the speed of the boat n still water and the speed of the current. Please help me I don't remember how to solve problems like this. xo Thanks

OpenStudy (anonymous):

Will you have two unknowns, boat speed and current speed. So you need two equations to solve the problem. The first equation for downstream. The second is for upstream. We are given that the round trip distance is 64 miles. So the distance downstream is half of 64 miles or 32 miles. The distance upstream is also half the round trip distance of 64 mile and it is 32 miles. Let b= the speed of the boat. Let c= the speed of the current. So the downstream equation would look like this: 4(b+c) = 32 The upstream equation would look like this: 8(b-c) = 32 So we have these equations: 4(b+c) = 32 8(b-c) = 32 Hopefully, you can finish solving from here.